OpenStudy (anonymous):

determine the stopping distances for a car with an initial speed of 95 kmhr-1 and human reaction time of 1 second for acceleration of: -4 ms-l and -8ms-1

OpenStudy (abhisar):

use the equation s=ut+\(\sf \frac{1}{2}t^2\)

OpenStudy (abhisar):

s=stopping distance u=initial velocity t=reaction time a=acceleration

OpenStudy (aum):

Use the equation \(v^2 = u^2 + 2as\) to compute the distance traveled by the car AFTER the brakes are applied. To this add the distance the car travels in one second which is the reaction time. That is, the person sees something and wants to apply the brakes to stop the car but it takes one second before the person applies the brakes.